# Time Weaver: A Conditional Time Series Generation Model

### v1

@@ Abstract
Imagine generating a city’s electricity demand pattern based on weather, the presence of an electric vehicle, and location, which could be used for capacity planning during a winter freeze. Such real-world time series are often enriched with paired heterogeneous contextual metadata (weather, location, etc.). Current approaches to time series generation often ignore this paired metadata, and its heterogeneity poses several practical challenges in adapting existing conditional generation approaches from the image, audio, and video domains to the time series domain. To address this gap, we introduce Time Weaver, a novel diffusion-based model that leverages the heterogeneous metadata in the form of categorical, continuous, and even time-variant variables to significantly improve time series generation. Additionally, we show that naive extensions of standard evaluation metrics from the image to the time series domain are insufficient. These metrics do not penalize conditional generation approaches for their poor specificity in reproducing the metadata-specific features in the generated time series. Thus, we innovate a novel evaluation metric that accurately captures the specificity of conditional generation and the realism of the generated time series. We show that Time Weaver outperforms state-of-the-art benchmarks, such as Generative Adversarial Networks (GANs), by up to 27% in downstream classification tasks on real-world energy, medical, air quality, and traffic data sets.
Machine Learning, ICML
[…]
@@ 1 Introduction
Despite potential advantages, current time series generation methods (Yoon et al., 2019; Jeha et al., 2021; Donahue et al., 2019) ignore the rich contextual metadata and, therefore, cannot be flexibly used to generate time series for specific real-world conditions. This is not due to a lack of data, as standard time series datasets have long come with paired metadata conditions. Instead, it is because today’s methods are incapable of handling diverse conditions.
@@
    Rich Metadata: Metadata can be categorical (e.g., whether a patient has a pacemaker), quantitative (e.g., age), or even a time series, such as anticipated precipitation. Any conditional generative model for time series should incorporate such a diverse mix of metadata conditions, as shown in Table 1. In contrast, image, video, and audio generation often deal with static text prompts.
@@
    Architectural Differences: In the image and audio domains, we have powerful feature extractors trained on internet-scale data (Radford et al., 2021; Wu* et al., 2023). These are vital building blocks for encoding conditions in image generation (Rombach et al., 2021). However, these models are non-existent in the time series domain due to the highly irregular nature of the time series datasets with respect to horizon lengths, number of channels, and the heterogeneity of the metadata.
@@
    Evaluation Metrics: Evaluating conditional generation approaches requires a metric that captures the specificity of the generated samples with respect to its paired metadata. In Fig. 4, we show how the existing metrics, such as the time series equivalent of the standard Frechet Inception Distance (FID) score, (Jeha et al., 2021), fail to capture this specificity and only measure how close the real and generated data distributions are. This is due to the fact that these metrics completely ignore the paired metadata in their evaluation.
@@
    We present Time Weaver (Fig. 1), a novel diffusion model for generating realistic multivariate time series conditioned on the paired metadata. We specifically innovate on the standard diffusion model architecture to process categorical and continuous metadata conditions.
@@
    We propose a new metric, the Joint Frechet Time Series Distance (J-FTSD), specifically designed to evaluate conditional time series data generation models. J-FTSD incorporates time series and metadata conditions using feature extractors trained with constrastive learning. In Sec. 6, we showcase J-FTSD’s ability to accurately rank approaches based on their ability to model conditional time series data distributions.
@@ 2 Background and Related Works
Figure 2: Time Weaver beats GAN s on all datasets for both J-FTSD and Train on Synthetic Test on Real (TSTR) metrics. We show percentage improvement of Time Weaver over state-of-the-art GAN models on four diverse datasets.
Generative Models in Time Series: Recently, Generative Adversarial Networks (GANs)(Donahue et al., 2019; Yoon et al., 2019; Li et al., 2022; Thambawita et al., 2021) have emerged as popular methods for time series data generation. However, these GAN-based approaches often struggle with unstable training and mode collapse (Chen, 2021). In response, Diffusion Models (DMs) (Sohl-Dickstein et al., 2015) have been introduced in the time series domain (Alcaraz & Strodthoff, 2023; Tashiro et al., 2021), offering more realistic data generation. DMs are a class of generative models that are shown to be state-of-the-art in a variety of domains, including image (Dhariwal & Nichol, 2021; Ho et al., 2020), speech (Chen et al., 2020; Kong et al., 2021), and video generation (Ho et al., 2022). DMs operate by defining a Markovian forward process $q$ by gradually adding noise to the clean data sample $x_{0}\sim\mathcal{X}$ where $\mathcal{X}$ is the data distribution to be learned. The forward process is predetermined by fixing a noise variance schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$, where $\beta_{t}\in\left[0,1\right]$ and $T$ is the total number of diffusion steps. The following equations describe the forward process:
@@
During training, a clean sample $x_{0}$ is transformed into $x_{t}$ using Eq. 2. Then, a neural network, $\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t)$, is trained to estimate the amount of noise added between $x_{t-1}$ and $x_{t}$ with the following loss function:
@@
Here, $t\sim\mathcal{U}(1,T)$ indicates that $t$ is sampled from a uniform distribution between 1 and $T$, $\epsilon$ is the noise added to $x_{t-1}$ to obtain $x_{t}$. In inference, we start from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathcal{N}(\mathbf{0},\mathbf{I})$ represents a zero mean, unit variance Gaussian distribution, and iteratively denoise using $\mathcal{\theta_{\textnormal{denoiser}}}$ to obtain a clean sample from the data distribution $\mathcal{X}$, i.e., $x_{T}\rightarrow x_{T-1},\ldots,x_{0}$. A detailed explanation of DMs is provided in App. A.
To extend to conditional DMs, the most commonly used approach is to keep the forward process the same as in Eq. 2 and add additional conditions $c$ to the reverse process. Minimizing $\left||\epsilon-\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t,c)\right\|^{2}_{2}$ in the loss function provided in Eq. 3 facilitates learning the conditional distribution. Conditional DMs are used in image, video (Saharia et al., 2021; Lugmayr et al., 2022; Rombach et al., 2021; Ramesh et al., 2022), and speech (Kong et al., 2021) generation. These models allow for diverse conditioning inputs, such as text, image, or even segmentation maps. However, these methods rely on image-focused tools like Convolutional Neural Networks (CNNs), which struggle to maintain essential time series characteristics, including long-range dependencies, as noted in (Gu et al., 2022). For time series data, models such as CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022) exist but are mainly limited to imputation tasks without substantial conditioning capabilities. Closest to our work, Alcaraz & Strodthoff (2023) attempts to incorporate ECG statements as metadata (only categorical) for ECG generation. However, this approach falls short as it does not consider heterogeneous metadata. Our method surpasses these limitations by effectively handling a broader range of metadata modalities, thus enabling more realistic time series data generation under varied heterogeneous conditions.
Metrics for Conditional Time Series Generation: Various metrics have been developed in the time series domain, focusing on the practical utility of the generated time series data. To this end, train on synthetic test on real (TSTR) metric (Jordon et al., 2018; Esteban et al., 2017) is used to assess the synthetic data’s ability to capture key features of the real dataset. TSTR metrics have been widely used to evaluate unconditional time series generation. Yoon et al. (2019) proposed the predictive score where synthetic time series data is used to train a forecaster, and forecast performance is evaluated on real time series data. More traditional approaches include average cosine similarity, Jensen distance (Li et al., 2022), and autocorrelation comparisons (Lin et al., 2020; Bahrpeyma et al., 2021). However, these heuristics often fail to fully capture the nuanced performance of conditional generative models.
@@
To evaluate image generation models, the FD is adjusted to the FID (Heusel et al., 2017). FID uses a feature extractor, the Inceptionv3 model (Szegedy et al., 2015), to transform images into embeddings, upon which the FD is calculated. Similar adaptations such as Frechet Video Distance (Unterthiner et al., 2018), Frechet ChemNet Distance (Preuer et al., 2018), and Context-FID (Jeha et al., 2021) exist for other domains, employing domain-specific feature extractors. However, these metrics are designed only to evaluate unconditional data generation since they only match the true data distribution marginalizing over all the conditions.
To evaluate conditional generation models, many metrics are proposed for categorical conditions (Murray, 2019; Huang et al., 2018; Benny et al., 2020; Liu et al., 2018; Miyato & Koyama, 2018). To create a more general metric, Soloveitchik et al. (2022) proposed the conditional FID (CFID) metric that works with continuous conditionals and calculates the conditional distributions of the generated and real data given the condition. In particular, DeVries et al. (2019) propose the Frechet Joint Distance (FJD), where the embeddings of the image and condition are obtained with different embedding functions and concatenated to create a joint embedding space. DeVries et al. (2019) consider conditions that are classes (image category), text descriptions (image captions), or images (for tasks like style transfer). However, in our case, the metadata could be any arbitrary combination of categorical, continuous conditions that might vary over time. Additionally, like other metrics considered in the literature, FJD is defined for image generation and does not consider the unique characteristics of time series data. In contrast, our proposed metric, J-FTSD, is specifically designed for evaluating time series data generation models conditioned on heterogeneous metadata.
@@ 3 Problem Formulation
Figure 3: Time Weaver architecture for incorporating metadata in the diffusion process: This figure shows the training process of Time Weaver model. Starting from the original sample $x_{0}$ (on the left), we gradually add noise through a forward process $q(x_{t}\mid x_{t-1})$ resulting in noisy samples $x_{t}$. In the reverse process, first, the categorical and continuous metadata are preprocessed with tokenizers $\theta_{\textnormal{token}}^{\textnormal{cat}}$ and $\theta_{\textnormal{token}}^{\textnormal{cont}}$ respectively. Then, we concatenate their output and process it through a self-attention layer $\theta_{\textnormal{condn}}$ to create the metadata embedding $z$. This embedding, along with the noisy sample $x_{t}$, is fed into the denoiser model $\mathcal{\theta_{\textnormal{denoiser}}}$. All the models are trained jointly to iteratively reconstruct a less noisy sample $x_{t-1}$. This denoising process is repeated until the original sample $x_{0}$ is reconstructed.
Consider a multivariate time series sample $x\in\mathbb{R}^{L\times F}$, where $L$ denotes the time series horizon and $F$ denotes the number of channels. Each sample $x$ is associated with metadata $c$, comprising categorical features $c_{\textnormal{cat}}\in\mathbb{N}^{L\times K_{\textnormal{cat}}}$ and continuous features $c_{\textnormal{cont}}\in\mathbb{R}^{L\times K_{\textnormal{cont}}}$. Here, $K_{\textnormal{cat}}$ and $K_{\textnormal{cont}}$ indicate the total numbers of categorical and continuous metadata features, respectively. These features are concatenated as $c=c_{\textnormal{cat}}\oplus c_{\textnormal{cont}}$, where $\oplus$ represents the vector concatenation operation. Thus, the metadata domain is defined as $\mathbb{N}^{L\times K_{\textnormal{cat}}}\times\mathbb{R}^{L\times K_{\textnormal{cont}}}$. Note that the domains of $c_{\textnormal{cat}}$ and $c_{\textnormal{cont}}$ allow time-varying metadata features.
Example: Consider generating time series data representing traffic volume on a highway ($F=1$) over a 96-hour period ($L=96$), using paired metadata. This metadata includes seven time-varying categorical features such as holidays (12 unique labels) and weather descriptions (11 unique labels), denoted by $K_{\textnormal{cat}}=7$. It also includes four time-varying continuous features like expected temperature and rain forecast, represented by $K_{\textnormal{cont}}=4$.
We denote the dataset $D_{x,c}=\{(x_{i},c_{i})\}_{i=1}^{n}$ consisting of $n$ independent and identically distributed (i.i.d) samples of time series data $x$ and paired metadata $c$, sampled from a joint distribution $p(x,c)$. Our objective is to develop a conditional generation model $G$, such that the samples generated by $G(c)$, distributionally match $p(x|c)$.
@@ 4 Conditional Time Series Generation using Time Weaver
Our approach, Time Weaver, is a diffusion-based conditional generation model. We choose DMs over GANs as we consider heterogenous metadata, i.e., the metadata can contain categorical, continuous, or even time-varying features. Previous works show that the conditional variants of GANs suffer from mode collapse when dealing with continuous conditions (Ding et al., 2020). Additionally, the proposed alternatives have not been tested in the time series domain for heterogeneous metadata. Our Time Weaver model consists of two parts - a denoiser backbone that generates data and a preprocessing module that processes the time-varying categorical and continuous metadata variables.
@@
Metadata Preprocessing: The preprocessing step involves handling the metadata $c=c_{\textnormal{cat}}\oplus c_{\textnormal{cont}}$. Here, $c_{\textnormal{cat}}\in\mathbb{N}^{L\times K_{\textnormal{cat}}}$ and $c_{\textnormal{cont}}\in\mathbb{R}^{L\times K_{\textnormal{cont}}}$ represent time-varying categorical and continuous metadata features respectively (see Sec. 3). To better incorporate these features from different modalities, we process them separately and then combine them with a self-attention layer.
@@
  • The categorical tokenizer $\theta_{\textnormal{token}}^{\textnormal{cat}}$ first converts each category in $c_{\textnormal{cat}}$ into one-hot encoding and then processes with fully connected (FC) layers to create categorical embedding $z_{\textnormal{cat}}\in\mathbb{R}^{L\times d_{\textnormal{cat}}}$. Similarly, the continuous tokenizer $\theta_{\textnormal{token}}^{\textnormal{cont}}$ also uses FC layers to encode continuous metadata $c_{\textnormal{cont}}$ into continuous embeddings $z_{\textnormal{cont}}\in\mathbb{R}^{L\times d_{\textnormal{cont}}}$. Using FC layers allows the model to learn the inherent correlation between the different metadata features within the categorical and continuous domains. Using FC layers is just a design choice, and more sophisticated layers can also be used.
@@
  • $z_{\textnormal{cat}}$ and $z_{\textnormal{cont}}$ are then concatenated and passed into a self-attention layer $\theta_{\textnormal{condn}}$ to generate the metadata embedding $z\in\mathbb{R}^{L\times d_{\textnormal{meta}}}$. The self-attention layer equips the generative model to capture the temporal relationship between the different metadata features.
@@
Here, $d_{\textnormal{cat}}$, $d_{\textnormal{cont}}$, and $d_{\textnormal{meta}}$ are design choices, and we refer the reader to App. D for further details.
@@
Denoiser: As the denoiser backbone for Time Weaver, we rely on two state-of-the-art architectures - CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022). The CSDI model uses feature and temporal self-attention layers to process sequential time series data, while SSSD uses structured state-space layers. Note that these denoiser models are designed for imputation and forecasting tasks, so they are designed to take unimputed and historical time series as respective inputs. We modify these denoisers into more flexible metadata-conditioned time series generators by augmenting them with preprocessing layers ($\theta_{\textnormal{token}}^{\textnormal{cat}}$, $\theta_{\textnormal{token}}^{\textnormal{cont}}$, and $\theta_{\textnormal{condn}}$). We refer the reader to App. D for details regarding architectural changes. We train the preprocessing layers $\theta_{\textnormal{condn}}$, $\theta_{\textnormal{token}}^{\textnormal{cont}}$, and $\theta_{\textnormal{token}}^{\textnormal{cat}}$, and the denoiser $\mathcal{\theta_{\textnormal{denoiser}}}$ jointly with the following loss:
@@
$\mathcal{L}_{(\mathcal{\theta_{\textnormal{denoiser}}},\theta_{\textnormal{condn}},\theta_{\textnormal{token}}^{\textnormal{cont}},\theta_{\textnormal{token}}^{\textnormal{cat}})}=$ $\mathbb{E}_{x,c\sim D_{x,c},\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I)},t\sim\mathcal{U}(1,T)}$
@@
$\left[\|\epsilon-\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t,z)\right\|^{2}_{2}],$ (5)
@@
where $z=\theta_{\textnormal{condn}}(\theta_{\textnormal{token}}^{\textnormal{cat}}(c_{\textnormal{cat}})\oplus\theta_{\textnormal{token}}^{\textnormal{cont}}(c_{\textnormal{cont}}))$, $D_{x,c}$ represents the dataset of time series and paired metadata sampled from the joint distribution $p(x,c)$, and $T$ is the total number of diffusion steps. As explained in Sec. 2, minimizing the loss in Eq. 5, allows Time Weaver to learn how to generate samples from the conditional distribution $p(x|c)$. During inference, we start from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively denoise (with metadata $c$ as input) for $T$ steps to generate $x_{0}\sim p(x|c)$. This process is depicted in Fig. 3.

### v2

@@ Abstract
Imagine generating a city’s electricity demand pattern based on weather, the presence of an electric vehicle, and location, which could be used for capacity planning during a winter freeze. Such real-world time series are often enriched with paired heterogeneous contextual metadata (e.g., weather and location). Current approaches to time series generation often ignore this paired metadata. Additionally, the heterogeneity in metadata poses several practical challenges in adapting existing conditional generation approaches from the image, audio, and video domains to the time series domain. To address this gap, we introduce Time Weaver, a novel diffusion-based model that leverages the heterogeneous metadata in the form of categorical, continuous, and even time-variant variables to significantly improve time series generation. Additionally, we show that naive extensions of standard evaluation metrics from the image to the time series domain are insufficient. These metrics do not penalize conditional generation approaches for their poor specificity in reproducing the metadata-specific features in the generated time series. Thus, we innovate a novel evaluation metric that accurately captures the specificity of conditional generation and the realism of the generated time series. We show that Time Weaver outperforms state-of-the-art benchmarks, such as Generative Adversarial Networks (GANs), by up to 30% in downstream classification tasks on real-world energy, medical, air quality, and traffic datasets.
Time series Generation, Diffusion Models, Conditional Generation Evaluation
@@ 1 Introduction
Despite potential advantages, current time series generation methods (Yoon et al., 2019; Jeha et al., 2021; Donahue et al., 2019) ignore the rich contextual metadata and are incapable of generating time series for specific real-world conditions. This is not due to a lack of data, as standard time series datasets have long come with paired metadata conditions. Instead, it is because today’s methods cannot handle diverse metadata conditions.
@@
    Rich Metadata: Metadata can be categorical (e.g., whether a patient has a pacemaker), quantitative (e.g., age), or even a time series, such as anticipated precipitation. Any conditional generative model for time series should incorporate such a diverse mix of metadata conditions (Table 1). In contrast, image, video, and audio generation often deal with static text prompts.
@@
    Architectural Differences: In the image and audio domains, we have powerful feature extractors trained on internet-scale data (Radford et al., 2021; Wu* et al., 2023). These are vital building blocks for encoding conditions in image generation (Rombach et al., 2021). However, these models are non-existent in the time series domain due to the irregular nature of the time series datasets with respect to horizon lengths, number of channels, and the heterogeneity of the metadata.
@@
    Evaluation Metrics: Evaluating conditional generation approaches requires a metric that captures the specificity of the generated samples with respect to their paired metadata. In Fig. 4, we show how the existing metrics, such as the time series equivalent of the standard Frechet Inception Distance (FID) score (Jeha et al., 2021), fail to capture this specificity and only measure how close the real and generated data distributions are. This is because these metrics completely ignore the paired metadata in their evaluation.
@@
    We present Time Weaver (Fig. 1), a novel diffusion model for generating realistic multivariate time series conditioned on metadata. We specifically innovate on the standard diffusion model architecture to process categorical and continuous metadata conditions.
@@
    We propose the Joint Frechet Time Series Distance (J-FTSD), specifically designed to evaluate conditional time series data generation models. J-FTSD incorporates time series and metadata conditions with feature extractors trained using a contrastive learning framework. In Sec. 6, we showcase J-FTSD’s ability to accurately rank approaches based on their ability to model conditional time series data distributions.
@@ 2 Background and Related Works
Figure 2: Time Weaver beats GAN s on all datasets on the Joint Frechet Time Series Distance (J-FTSD) and Train on Synthetic Test on Real (TSTR) metrics. J-FTSD indicates the distributional similarity between the generated and real time series datasets. Lower values of J-FTSD indicate that both generated and real time series distributions are closer. TSTR indicates the performance of a downstream task model trained on generated time series data and evaluated on real time series data. Higher values of TSTR indicate higher quality of the generated time series data. We show the percentage improvement of Time Weaver over state-of-the-art GAN models on four diverse datasets.
Generative Models in Time Series: Recently, Generative Adversarial Networks (GANs) (Donahue et al., 2019; Yoon et al., 2019; Li et al., 2022; Thambawita et al., 2021) have emerged as popular methods for time series data generation. However, these GAN-based approaches often struggle with unstable training and mode collapse (Chen, 2021). In response, Diffusion Models (DMs) (Sohl-Dickstein et al., 2015) have been introduced in the time series domain (Alcaraz & Strodthoff, 2023; Tashiro et al., 2021), offering more realistic data generation. DMs are a class of generative models that are state-of-the-art in a variety of domains, including image (Dhariwal & Nichol, 2021; Ho et al., 2020), speech (Chen et al., 2020; Kong et al., 2021), and video generation (Ho et al., 2022). DMs operate by defining a Markovian forward process $q$. The forward process gradually adds noise to a clean data sample $x_{0}\sim\mathcal{X}$, where $\mathcal{X}$ is the data distribution to be learned. The forward process is predetermined by fixing a noise variance schedule $\left\{\beta_{1},\ldots,\beta_{T}\right\}$, where $\beta_{t}\in\left[0,1\right]$ and $T$ is the total number of diffusion steps. The following equations describe the forward process:
@@
Here, $\mathcal{N}(\mu,\Sigma)$ represents a Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$. During training, a clean sample $x_{0}$ is transformed into $x_{t}$ using Eq. 2. Then, a neural network, $\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t)$, is trained to estimate the amount of noise added between $x_{t-1}$ and $x_{t}$ with the following loss function:
@@
Here, $t\sim\mathcal{U}(1,T)$ indicates that $t$ is sampled from a uniform distribution between 1 and $T$, and $\epsilon$ is the noise added to $x_{t-1}$ to obtain $x_{t}$. In inference, we start from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathcal{N}(\mathbf{0},\mathbf{I})$ represents a zero mean, unit variance Gaussian distribution, and iteratively denoise using $\mathcal{\theta_{\textnormal{denoiser}}}$ to obtain a clean sample from the data distribution $\mathcal{X}$, i.e., $x_{T}\rightarrow x_{T-1},\ldots,x_{0}$. A detailed explanation of DMs is provided in Section A.1.
For conditional DMs, the most commonly used approach is to keep the forward process the same as in Eq. 2, and add additional conditions $c$ to the reverse process. Minimizing $\left\|\epsilon-\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t,c)\right\|^{2}_{2}$ in the loss function provided in Eq. 3 facilitates learning the conditional distribution. Conditional DMs are used in image, video (Saharia et al., 2021; Lugmayr et al., 2022; Rombach et al., 2021; Ramesh et al., 2022), and speech (Kong et al., 2021) generation. These models allow for diverse conditioning inputs, like text, image, or even segmentation maps. However, these methods rely on image-focused tools like Convolutional Neural Networks (CNNs), which struggle to maintain essential time series characteristics such as long-range dependencies, as noted in (Gu et al., 2022). For time series data, models such as CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022) exist but are mainly limited to imputation tasks without substantial conditioning capabilities. Closest to our work, Alcaraz & Strodthoff (2023) attempt to incorporate ECG statements as metadata (only categorical) for ECG generation. However, this approach falls short as it does not consider heterogeneous metadata. Our method surpasses these limitations by effectively handling a broader range of metadata modalities, thus enabling more realistic time series data generation under varied heterogeneous conditions.
Metrics for Conditional Time Series Generation: Various metrics have been developed in the time series domain, focusing on the practical utility of the generated time series data. To this end, the Train on Synthetic Test on Real (TSTR) metric (Jordon et al., 2018; Esteban et al., 2017) is used to assess the ability of synthetic data to capture key features of the real dataset. TSTR metrics have been widely used to evaluate unconditional time series generation. Yoon et al. (2019) proposed the predictive score where synthetic time series data is used to train a forecaster, and the forecaster’s performance is evaluated on real time series data. More traditional approaches include average cosine similarity, Jensen distance (Li et al., 2022), and autocorrelation comparisons (Lin et al., 2020; Bahrpeyma et al., 2021). However, these heuristics often fail to fully capture the nuanced performance of conditional generative models.
@@
To evaluate image generation models, the FD is adjusted to the Frechet Inception Distance (FID) (Heusel et al., 2017). FID uses a feature extractor, the Inception-v3 model (Szegedy et al., 2015), to transform images into embeddings upon which the FD is calculated. Similar adaptations such as the Frechet Video Distance (Unterthiner et al., 2018), Frechet ChemNet Distance (Preuer et al., 2018), and Context-FID (Jeha et al., 2021) exist for other domains, employing domain-specific feature extractors. However, these metrics are designed only to evaluate unconditional data generation since they only match the true data distribution marginalizing over all the conditions.
To evaluate conditional generation models, many metrics are proposed for categorical conditions (Murray, 2019; Huang et al., 2018; Benny et al., 2020; Liu et al., 2018; Miyato & Koyama, 2018). To create a more general metric, Soloveitchik et al. (2022) proposed the conditional FID (CFID) metric that works with continuous conditionals and calculates the conditional distributions of the generated and real data given the condition. In particular, DeVries et al. (2019) propose the Frechet Joint Distance (FJD), where the embeddings of the image and condition are obtained with different embedding functions and concatenated to create a joint embedding space. DeVries et al. (2019) consider conditions that are classes (image category), text descriptions (image captions), or images (for tasks like style transfer). However, in our case, the metadata could be any arbitrary combination of categorical, continuous, and time-varying conditions. Additionally, like other metrics considered in the literature, FJD is defined for image generation and does not consider the unique characteristics of time series data. In contrast, our proposed J-FTSD metric is specifically designed to evaluate time series data generation models conditioned on heterogeneous metadata.
@@ 3 Problem Formulation
Figure 3: Time Weaver architecture for incorporating metadata in the diffusion process: This figure shows the training and inference processes of Time Weaver. For training, we start from the original sample $x_{0}$ (on the left) and gradually add noise through a forward process $q(x_{t}\mid x_{t-1})$, resulting in noisy samples $x_{t}$. The denoiser, $\mathcal{\theta_{\textnormal{denoiser}}}$, is trained to estimate the amount of noise added to obtain $x_{t}$ from $x_{t-1}$. During inference, the categorical and continuous metadata are first preprocessed with tokenizers $\theta_{\textnormal{token}}^{\textnormal{cat}}$ and $\theta_{\textnormal{token}}^{\textnormal{cont}}$, respectively. Then, we concatenate their output and process it through a self-attention layer $\theta_{\textnormal{condn}}$ to create the metadata embedding $z$. This embedding is fed into $\mathcal{\theta_{\textnormal{denoiser}}}$ with the noisy sample $x_{t}$ to obtain $x_{t-1}$. The denoising process is repeated for $T$ diffusion steps to obtain a clean sample similar to $x_{0}$.
Consider a multivariate time series sample $x\in\mathbb{R}^{L\times F}$, where $L$ denotes the time series horizon, and $F$ denotes the number of channels. Each sample $x$ is associated with metadata $c$, comprising categorical features $c_{\textnormal{cat}}\in\mathbb{N}^{L\times K_{\textnormal{cat}}}$ and continuous features $c_{\textnormal{cont}}\in\mathbb{R}^{L\times K_{\textnormal{cont}}}$. Here, $K_{\textnormal{cat}}$ and $K_{\textnormal{cont}}$ indicate the total numbers of categorical and continuous metadata features, respectively. These features are concatenated as $c=c_{\textnormal{cat}}\oplus c_{\textnormal{cont}}$, where $\oplus$ represents the vector concatenation operation. Thus, the metadata domain is defined as $\mathbb{N}^{L\times K_{\textnormal{cat}}}\times\mathbb{R}^{L\times K_{\textnormal{cont}}}$. Note that the domains of $c_{\textnormal{cat}}$ and $c_{\textnormal{cont}}$ allow time-varying metadata features.
Example: Consider generating time series data representing traffic volume on a highway ($F=1$) over 96 hours ($L=96$) using paired metadata. This metadata includes seven time-varying categorical features like holidays (12 unique labels) and weather descriptions (11 unique labels), denoted by $K_{\textnormal{cat}}=7$. It also includes four time-varying continuous features like expected temperature and rain forecast, represented by $K_{\textnormal{cont}}=4$.
We denote the dataset $D_{x,c}=\{(x_{i},c_{i})\}_{i=1}^{n}$ consisting of $n$ independent and identically distributed (i.i.d) samples of time series data $x$ and paired metadata $c$, sampled from a joint distribution $p(x,c)$. Our objective is to develop a conditional generation model $G$ such that the samples generated by $G(c)$ distributionally match $p(x|c)$.
@@ 4 Conditional Time Series Generation using Time Weaver
Our approach, Time Weaver, is a diffusion-based conditional generation model. We choose DMs over GANs as we consider heterogeneous metadata, i.e., the metadata can contain categorical, continuous, or even time-varying features. Previous works show that the conditional variants of GANs suffer from mode collapse when dealing with continuous conditions (Ding et al., 2020). Additionally, the proposed alternatives have not been tested in the time series domain for heterogeneous metadata. Our Time Weaver model consists of two parts - a denoiser backbone that generates data and a preprocessing module that processes the time-varying categorical and continuous metadata features.
@@
Metadata Preprocessing: The preprocessing step involves handling the metadata $c=c_{\textnormal{cat}}\oplus c_{\textnormal{cont}}$. Here, $c_{\textnormal{cat}}\in\mathbb{N}^{L\times K_{\textnormal{cat}}}$ and $c_{\textnormal{cont}}\in\mathbb{R}^{L\times K_{\textnormal{cont}}}$ represent time-varying categorical and continuous metadata features, respectively (see Sec. 3). To better incorporate these features from different modalities, we process them separately and then combine them with a self-attention layer.
@@
  • The categorical tokenizer $\theta_{\textnormal{token}}^{\textnormal{cat}}$ first converts each category in $c_{\textnormal{cat}}$ into one-hot encoding and then processes with fully connected (FC) layers to create the categorical embedding $z_{\textnormal{cat}}\in\mathbb{R}^{L\times d_{\textnormal{cat}}}$. Similarly, the continuous tokenizer $\theta_{\textnormal{token}}^{\textnormal{cont}}$ also uses FC layers to encode continuous metadata $c_{\textnormal{cont}}$ into the continuous embedding $z_{\textnormal{cont}}\in\mathbb{R}^{L\times d_{\textnormal{cont}}}$. These FC layers learn the correlation between metadata features within the categorical and continuous domains. Using FC layers is just a design choice; more sophisticated layers can also be used.
@@
  • $z_{\textnormal{cat}}$ and $z_{\textnormal{cont}}$ are then concatenated and passed through a self-attention layer $\theta_{\textnormal{condn}}$ to generate the metadata embedding $z\in\mathbb{R}^{L\times d_{\textnormal{meta}}}$. The self-attention layer equips the generative model to capture the temporal relationship between different metadata features.
@@
Here, $d_{\textnormal{cat}}$, $d_{\textnormal{cont}}$, and $d_{\textnormal{meta}}$ are design choices, and we refer the reader to Section A.4 for further details.
@@
Denoiser: As the denoiser backbone for Time Weaver, we rely on two state-of-the-art architectures - CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022). The CSDI model uses feature and temporal self-attention layers to process sequential time series data, while SSSD uses structured state-space layers. Note that these denoisers are designed for imputation and forecasting tasks. So, they are designed to take unimputed and historical time series as respective inputs. We modify these denoisers into more flexible metadata-conditioned time series generators by augmenting them with preprocessing layers ($\theta_{\textnormal{token}}^{\textnormal{cat}}$, $\theta_{\textnormal{token}}^{\textnormal{cont}}$, and $\theta_{\textnormal{condn}}$). We refer the reader to Section A.4 for details regarding architectural changes. We train the preprocessing layers and the denoiser $\mathcal{\theta_{\textnormal{denoiser}}}$ jointly with the following loss:
@@
$\mathcal{L}_{(\mathcal{\theta_{\textnormal{denoiser}}},\theta_{\textnormal{condn}},\theta_{\textnormal{token}}^{\textnormal{cont}},\theta_{\textnormal{token}}^{\textnormal{cat}})}=$ $\mathbb{E}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x,c\sim D_{x,c}},\epsilon\sim\mathcal{N}(\textbf{0},\textbf{I)},t\sim\mathcal{U}(1,T)}$
@@
$\left[\|\epsilon-\mathcal{\theta_{\textnormal{denoiser}}}(x_{t},t,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z})\right\|^{2}_{2}],$ (5)
@@
where $z=\theta_{\textnormal{condn}}(\theta_{\textnormal{token}}^{\textnormal{cat}}(c_{\textnormal{cat}})\oplus\theta_{\textnormal{token}}^{\textnormal{cont}}(c_{\textnormal{cont}}))$, $D_{x,c}$ represents the dataset of time series and paired metadata sampled from the joint distribution $p(x,c)$, and $T$ is the total number of diffusion steps. As explained in Sec. 2, minimizing the loss in Eq. 5 allows Time Weaver to learn how to generate samples from the conditional distribution $p(x|c)$. During inference, we start from $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively denoise (with metadata $c$ as input) for $T$ steps to generate $x_{0}\sim p(x|c)$. This process is depicted in Fig. 3.